4x - 3y = 10
y - 20 = -6x
Solve by addition method

Answer by checkley77(12844)

4x - 3y = 10
y - 20 = -6x multiply this equation by 3 & add
3y-60=-18x
-3y-10=-4x now add.
------------------------------
-70=-22x
x=-70/-22
x=35/11 answer.
y-20=-6*35/11
y-20=-210/11
y=-210/11+20
y=(-210+20*11)/11
y=(-210+220)/11
y=10/11 answer.
Proof:
4*35/11-3*10/11=10
140/11-30/11=10
110/11=10
10=10